d dir Q1] [20 points] Below, p(D) will denote a polynomial in differential operator D= For the given p(D), find the general sb) (5 points) p(D) = D(D-4)(D+3).c) (5 points) p(D) = (D2 + 3)?d) (5 points) p(D) = D2 – D+4
Answer #1

P(D) = D470+6 2 Auxilary equn - 3²+70+6=0 Dtod+D+6=0 (6+6) (D+1)=0 -6,-) (Real & distinct root D General solun -62 y = C, et2 PCD) - D - Dt 4 A-E 2_Dt 4 =0 D =-6 Į J6?_que a=1, 6=-1, (= 4 2a tit J-16 1 tsisi 2 don roots STS Complex D -100 זן sinuts